An apartment building has 20 two-bedroom apartments and 8 three-bedroom apartments. If 4 apartments are chosen at random to be redecorated, what are the probabilities that
A) All will be two-bedroom apartments?
B) 2 will be two-bedroom apartments and 2 will be three-bedroom apartments?

Answer by greenestamps(13334)



A) choose 4 of the 28 apartments, choosing 4 of the 20 2-bedroom apartments and 0 of the 8 3-bedroom apartments:

((20C4)*(8C0))/(28C4) = (4845*1)/20475 = 4845/20475 = 323/1365 = 0.2366 to 4 decimal places.

B) choose 4 of the 28 apartments, choosing 2 of the 20 2-bedroom apartments and 2 of the 8 3-bedroom apartments:

((20C2)*(8C2))/(28C4) = (190*28)/20475 = 5320/20475 = 152/585 = 0.2598 to 4 decimal places